OpenStudy (anonymous):

If all three dimensions of a rectangular prism double, then the volume will be three times larger.

OpenStudy (anonymous):

volume of a prism is V1 = lwh now if all dimensions are doubled, then the new volume will be V2 = (2l)(2w)(2h) = 2*2*2*lwh = 8*V1 now compare the original volume, V1, to the new volume, V2.